OpenStudy (anonymous):

Use the right triangle and values to find the missing length. Round to the nearest tenth. A = 2, B = 4, C = ?

OpenStudy (anonymous):

Use the pythagorean theorem x^2+b^2=c^2 2^2+4^2=c^2 4+16=c^2 20=c^2 Find the square root of each Square root of 20=C

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

I looked at the multiple answers that they have and 20 is not in them. it is : 1. 3 2. 2.8 3. 3.5 4. 4.5

OpenStudy (anonymous):

I said it was the square root of 20=4.47 which would be letter 4

OpenStudy (anonymous):

My bad , I didn't see that you wrote that. I thought it said that 20 was the answer. totally my fault but thanks for all your help !

OpenStudy (anonymous):

Haha it's alright :)